# Offline Clustering of Linear Bandits: Unlocking the Power of Clusters in Data-Limited Environments

### v1

@@ 7 Experiments
Experimental Setup. We consider $U=1k$ users, partitioned into $J=10$ clusters, with datasets ranging from $5k$ to $1M$ points, equally split between offline learning and evaluation. The reward model is $r_{u}^{i}=\langle\bm{\theta}_{u},\bm{a}_{u}^{i}\rangle+\epsilon_{u}^{i}$, where $\bm{a}_{u}^{i}$ and $\bm{\theta}_{u}$ are item and user features, and $\epsilon_{u}^{i}$ is 1-sub-Gaussian noise. We benchmark against LinUCB_IND, CLUB, an enhanced DBSCAN variant (DBSCAN_Improve), and an improved variant of XMeans (Pelleg et al., 2000) (XMeans_Improve). We use both a synthetic dataset, consisting of $d=20$-dimensional Gaussian vectors, and two real-world datasets (Yelp [^1] https://business.yelp.com/data/resources/open-dataset/. and MovieLens Harper and Konstan (2015)), sampled from the top 1,000 users and items. User preference vectors are extracted using Singular Value Decomposition (SVD). Offline data is generated using random or LinUCB-based item selection. More details can be found in Appendix C.
@@
Performance with Insufficient Dataset. Figure 2 shows the average suboptimality gap under different user distributions (equal and semi-random) and dataset sizes ($|\mathcal{D}|$). Each subplot represents a unique user distribution and dataset. For Off-C2LUB, $\hat{\gamma}$ is optimized in an equal distribution scenario and applied to the semi-random distribution. When $|\mathcal{D}|\in[20k,100k]$, Off-C2LUB_Optimistic (Pessimistic) significantly reduces the suboptimality gap. On the synthetic dataset, it outperforms Off-CLUB by 33.0% (31.7%) and other baselines by at least 84.7% (85.9%). On Yelp, it reduces the gap by 55.9% (58.8%) over Off-CLUB and at least 73.1% (74.8%) over baselines. On MovieLens, it reduces the gap by 41.3% (44.9%) over Off-CLUB and at least 66.0% (68.1%) over baselines.
Performance with Sufficient Dataset. Table 3 shows that Off-CLUB performs better as $|\mathcal{D}|$ increases and it is more likely to have sufficient data. Furthermore, as $|\mathcal{D}|$ increases, both Off-CLUB and Off-C2LUB consistently maintain superior performance, achieving the lowest suboptimality gaps.
Improvement with Estimated $\hat{\gamma}$. We evaluate various strategies under insufficient data using synthetic and Yelp datasets (Figure 3), with $|\mathcal{D}|=30k$; additional results are in Appendix C. Solid lines indicate methods that use $\hat{\gamma}$, dashed lines those that do not. On the synthetic dataset, relative suboptimality gaps compared to the optimal $\hat{\gamma}$ of Off-C2LUB are: DBSCAN (5.90×5.90$\times$5.90 ×), CLUB (2.55×2.$5$), Off-C2LUB_Pessimistic (1.50×1.50$\times$1.50 ×), and Off-C2LUB_Optimistic (1.22×1.$2$). Similar trends hold for Yelp: DBSCAN (3.10×3.10$\times$3.10 ×), CLUB (2.26×2.26$\times$2.26 ×), Pessimistic (1.41×1.41$\times$1.41 ×), and Optimistic (1.24×1.24$\times$1.24 ×).
Figure 2: Suboptimality Gap under Varying User Distributions with Limited Data
@@
Figure 3: Suboptimality Gap under Varying $\hat{\gamma}$
Table 2: Effect of Data Size on Suboptimality
@@
DBSCAN_Improve 0.130149 0.058302 0.024758 0.011941 0.006520
XMeans_Improve 0.129364 0.058108 0.024791 0.011955 0.006536
CLUB 0.013153 0.000019 0.000012 0.000009 0.000007
Off-CLUB 0.000038 0.000018 0.000012 0.000009 0.000007
LinUCB 0.140748 0.140993 0.140925 0.140898 0.140844
LinUCB_IND 0.144851 0.066494 0.028189 0.013586 0.007561
SCLUB 0.180909 0.131105 0.050664 0.051058 0.051328
Off-C2LUB_Pessimistic 0.000037 0.000019 0.000013 0.000011 0.000008
Off-C2LUB_Optimistic 0.000041 0.000018 0.000012 0.000009 0.000007
@@ 8 Conclusion
Appendix A Detailed Proofs
A.1 Proof of Theorem 4.2
To prove Theorem 4.2, we first introduce several key notations and supporting lemmas, which we present and prove sequentially.
Table 4: Notations Defined in Proofs
Notation Description
$V_{\hat{\gamma}}(u)$ The set consisting of user $u$ and its neighbors in the graph $\mathcal{G}_{\hat{\gamma}}$.
$R_{\hat{\gamma}}(u)$ Users in $V_{\hat{\gamma}}(u)$ that share the same preference vector as user $u$.
$W_{\hat{\gamma}}(u)$ Users in $V_{\hat{\gamma}}(u)$ that do not share the same preference vector as user $u$.
$W_{\hat{\gamma}}^{j}(u)$ Users in $W_{\hat{\gamma}}(u)$ who belong to cluster $V(j)$.
$J_{\hat{\gamma}}(u)$ Indices of distinct clusters represented in $W_{\hat{\gamma}}(u)$.
To characterize the contributions of different types of neighbors in the estimation process, we define the weighted fraction of samples from $R_{\hat{\gamma}}(u)$ relative to $V_{\hat{\gamma}}(u)$ as:
$\eta_{R_{\hat{\gamma}}(u)}=\frac{|R_{\hat{\gamma}}(u)|\lambda+N^{R_{\hat{\gamma}}(u)}}{|V_{\hat{\gamma}}(u)|\lambda+N^{V_{\hat{\gamma}}(u)}}.$
Here, $\lambda$ is the regularization parameter used in constructing each user’s matrix $M_{u}$. Analogous definitions apply to $\eta_{W_{\hat{\gamma}}(u)}$ and $\eta_{W_{\hat{\gamma}}^{j}(u)}$, and by construction, we have:
$\eta_{R_{\hat{\gamma}}(u)}+\eta_{W_{\hat{\gamma}}(u)}=1,$
since $V_{\hat{\gamma}}(u)=R_{\hat{\gamma}}(u)\cup W_{\hat{\gamma}}(u)$ and the two sets are disjoint.
Lemma A.1.
With the notations defined in Table 4, the suboptimality gap of Off-C2LUB satisfies:
@@ Appendix A Detailed Proofs
$\textnormal{SubOpt}(u_{\textnormal{test}},\textnormal{Off-C}^{2}\textnormal{LUB},A_{\textnormal{test}})\leq\frac{(2\sqrt{2}+1)\beta(V_{\hat{\gamma}}(u_{\text{test}}),\delta)}{\sqrt{\lambda_{a}N^{V_{\hat{\gamma}}(u_{\text{test}})}}}+\frac{4\lambda U}{\lambda_{a}N^{V_{\hat{\gamma}}(u_{\text{test}})}}+\frac{4\eta_{W_{\hat{\gamma}}(u_{\text{test}})}}{\lambda_{a}}\cdot\hat{\gamma}$
with probability at least $1-2\delta$, where
$\beta(\mathcal{S},\delta)=\sqrt{d\log\left(1+\frac{N^{\mathcal{S}}}{|\mathcal{S}|\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{|\mathcal{S}|\lambda}$
for any set $\mathcal{S}$.
Lemma A.2.
The bound in Lemma A.1 can be simplified as:
[…]
Lemma A.3.
For the notations and inputs in Theorem 4.2, the following containment relations hold:
$\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})\subseteq R_{\hat{\gamma}}(u_{\text{test}}),\quad\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})\supseteq W_{\hat{\gamma}}(u_{\text{test}}),\quad\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}\geq\eta_{W_{\hat{\gamma}}(u_{\text{test}})}.$
Proof of Lemma A.1.
@@
Let $\beta(\mathcal{S},\delta)=\sqrt{d\log\left(1+\frac{\sum_{u\in\mathcal{S}}N_{u}}{|\mathcal{S}|\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{|\mathcal{S}|\lambda}$ for any set $\mathcal{S}$. We begin by defining the following confidence event for each user $u$:
@@
By Lemma B.1, we have $\mathbb{P}[\mathcal{F}_{v}^{\prime}(\delta)]\geq 1-\frac{\delta}{2U}$ for each $v\in\mathcal{U}$. Therefore, applying the union bound over all users yields:
@@
Furthermore, for both $R_{\hat{\gamma}}(u)$ and $W_{\hat{\gamma}}(u)$, the clustering condition in Equation 2 requires that a user $v$ can be included in either set only if both $N_{u}$ and $N_{v}$ are at least $\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right)$. Hence, the case where $N_{u}<\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right)$ is trivial since the problem can be simplified to an offline linear contextual bandit problem in [Li et al., 2022].
For the non-trivial case where $N_{u}\geq\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right)$, we apply Lemma B.2 and the union bound to obtain that for any $v\in R_{\hat{\gamma}}(u)\cup W_{\hat{\gamma}}(u)$, with probability at least $1-\frac{3}{2}\delta$, the following inequality holds:
$\left\|\bm{\theta}_{v}-\hat{\bm{\theta}}_{v}\right\|_{2}\leq\frac{\left\|\bm{\theta}_{v}-\hat{\bm{\theta}}_{v}\right\|_{M_{v}}}{\sqrt{\lambda_{\min}(M_{v})}}\leq\frac{\beta(\{v\},\delta)}{\sqrt{\lambda+\lambda_{a}N_{v}/2}}\leq\text{CI}_{v}.$

### v2

@@ 6 Experiments
Figure 2: Comparisons of our algorithms with baselines under different user distributions and datasets. Off-C2LUB and its variants consistently outperform Off-CLUB and other baselines.
Experimental Setup. We evaluate on both synthetic and real-world datasets, including Yelp (yel, 2023) and MovieLens (Harper and Konstan, 2015). We fix $U=1\text{k}$ users with preference dimension $d=20$, and consider dataset sizes from $5\text{k}$$100\text{k}$ (insufficient data) and $0.2\text{M}$$1\text{M}$ (sufficient data), splitting each dataset evenly for training and evaluation. Synthetic users are divided into $J=10$ clusters under two sampling schemes: Equal Distribution (uniform over users) and Semi-Random Distribution (uniform within clusters, non-uniform across clusters). For real-world data, we keep the top 1,000 users and items by rating count and derive preference vectors via SVD (Li et al., 2019). All results are averaged over 10 random runs. Baselines include bandit methods (LinUCB (Abbasi-Yadkori et al., 2011), CLUB (Gentile et al., 2014), SCLUB (Li et al., 2019)) and clustering methods (DBSCAN (Schubert et al., 2017), XMeans (Pelleg et al., 2000), ARMUL (Duan and Wang, 2023)), adapted to the offline setting. Offline data is generated either by random selection (Figure 2) or LinUCB-based item selection (Appendix E.3). Further details are given in Appendices E.1E.3.
@@
Performance with Insufficient Dataset. To illustrate the performance under insufficient data, we report the average suboptimality gap across user distributions and dataset sizes (Figure 2). For Off-C2LUB, $\hat{\gamma}$ is optimized under the equal distribution scenario and applied to the semi-random distribution. With $|\mathcal{D}|\in[20k,100k]$, both Off-C2LUB variants significantly reduce the suboptimality gap. On the synthetic dataset, Off-C2LUB_Overestimation improves over Off-CLUB by 44.2% and over other baselines by at least 77.5%, while Off-C2LUB_Underestimation achieves 34.0% and 73.3%, respectively. On Yelp, the corresponding improvements are 19.8% and 73.9% for Overestimation, and 17.9% and 73.3% for Underestimation. On MovieLens, they are 8.6% and 67.9% for Overestimation, and 9.3% and 68.0% for Underestimation. Overall, Off-C2LUB achieves the best performance, while Off-CLUB, though less effective, still outperforms classical clustering baselines.
Performance with Sufficient Data. Table 3 reports the performance of different algorithms under sufficient data. The bold values mark the optimal suboptimal gaps (rounded to six decimals) for each dataset size. Our proposed Off-CLUB and Off-C2LUB consistently outperform all other baselines in this regime, with the $\hat{\gamma}$ selection strategies for Off-C2LUB also guarantee near optimal performances. This experiment shows that with sufficient data, both Off-CLUB and Off-C2LUB approach to near-zero suboptimality as more offline data is provided, outperforming existing methods.
Improvement with Estimated $\hat{\gamma}$. We evaluate various strategies under insufficient data using synthetic and Yelp datasets (Figure 3) with $|\mathcal{D}|=30k$. The results highlight that the suboptimality gap of Off-C2LUB increases when $\hat{\gamma}$ is set either too small or too large, which aligns with our theoretical analysis of the noise–bias tradeoff in Lemma˜4.2, Theorem˜5.2, and Figure˜1. On both the synthetic and Yelp datasets, the proposed overestimation and underestimation strategies achieve near-optimal performance. Additional experiments are provided in Appendix E, where we further examine the robustness of our methods to varying dataset sizes and datasets.
@@
Figure 3: Suboptimality under varying $\hat{\gamma}$, showing Off-C2LUB is near-optimal with our overestimation or underestimation strategies.
Table 3: Comparisons under sufficient data show Off-CLUB and Off-C2LUB reach near-zero suboptimality, outperforming baselines.
@@
DBSCAN_Improve 0.130230 0.058202 0.024840 0.011893 0.006512
XMeans_Improve 0.130088 0.058248 0.024820 0.011908 0.007534
CLUB 0.144795 0.066638 0.028110 0.013647 0.007557
Off-CLUB 0.000037 0.000018 0.000013 0.000009 0.000007
LinUCB_IND 0.144894 0.066417 0.028238 0.013664 0.007580
SCLUB 0.001633 0.000447 0.000165 0.000095 0.000057
ARMUL 0.146900 0.109480 0.078815 0.055882 0.051504
Off-C2LUB 0.000070 0.000020 0.000012 0.000009 0.000007
Off-C2LUB_Underestimation 0.000037 0.000019 0.000013 0.000010 0.000008
Off-C2LUB_Overestimation 0.000042 0.000018 0.000012 0.000009 0.000007
@@ 7 Conclusion
Appendix A Additional Discussions on Action Regularity Assumptions
In this appendix, we provide a detailed discussion of ˜3.1. This assumption is standard in the clustering of bandits literature [Gentile et al., 2014, Li and Zhang, 2018, Li et al., 2019, Liu et al., 2022, Wang et al., 2023a, Dai et al., 2024a, Li et al., 2025, Wang et al., 2025], although the precise formulation varies some across works. The version we adopt in this paper—where each offline action is drawn from a candidate set $\mathcal{S}_{u}^{i}$ of bounded size $S$ and the algorithm is assumed to know the smoothed regularity parameter $\tilde{\lambda}_{a}$—follows the most recent line of research [Wang et al., 2023a, Dai et al., 2024a, Wang et al., 2025]. To highlight the differences, we first introduce an alternative assumption that can remove the finite candidate set requirement or the need for $\tilde{\lambda}_{a}$, at the cost of stronger independence or bounded-variance conditions. We then discuss how our analysis extends to cases where no regularity assumption is imposed.
Assumption A.1 (Action Regularity (Alternative Form)).
Let $\rho$ be a distribution over $\{\bm{a}\in\mathbb{R}^{d}:\|\bm{a}\|_{2}\leq 1\}$ such that $\mathbb{E}_{\bm{a}\sim\rho}[\bm{a}\bm{a}^{\top}]$ is full rank with minimum eigenvalue $\lambda_{a}>0$. Each action $\bm{a}_{u}^{i}$ in dataset $\mathcal{D}_{u}$ is either: (i) independently and identically drawn from $\rho$, or (ii) selected from a candidate set $\mathcal{S}_{u}^{i}$, where actions in $\mathcal{S}_{u}^{i}$ are independently drawn from $\rho$ and, for any unit vector $\bm{\theta}\in\mathbb{R}^{d}$, the random variable $(\bm{\theta}^{\top}\bm{a})^{2}$ has sub-Gaussian tails with variance $\sigma^{2}\leq\lambda_{a}^{2}/(8\log(4|\mathcal{S}_{u}^{i}|))$. We assume $\lambda_{a}$ is known to the algorithm.
One crucial distinction between ˜A.1 and ˜3.1 lies in the prior knowledge required: the former requires $\lambda_{a}$, whereas the latter uses the smoothed regularity parameter $\tilde{\lambda}_{a}$. When ˜3.1 does not hold but ˜A.1 does, all our theoretical claims (such as Theorem˜4.4 and Theorem˜5.2) remain valid after replacing every occurrence of $\tilde{\lambda}_{a}$ with $\lambda_{a}$ in the pseudo-code and analysis of Algorithms 1 and 2, according to Lemma˜D.2 and D.3.
Interpretation of the Alternative Form.
˜A.1 accommodates two ways of generating offline data:
(1) The i.i.d. action model. Each action is drawn i.i.d. from $\rho$, as motivated by the online exploration phase proposed by algorithms in recent online clustering of bandits work [Li et al., 2025], and widely adopted in offline RL/bandits literature [Rashidinejad et al., 2021, Liu et al., 2025]. This setting is appropriate when offline data are sampled randomly from a continuous or otherwise infinite action space. However, it does not capture cases where actions are selected by a logging policy (e.g., LinUCB [Abbasi-Yadkori et al., 2011]) which is common in real-world offline datasets and introduces dependencies, since a logging policy may violate the independency requirement.
(2) Dependent actions with bounded variance. Following early clustering works [Gentile et al., 2014, Li and Zhang, 2018, Li et al., 2019, Liu et al., 2022], this model allows dependencies among actions but requires an additional bounded-variance condition. This assumption still relies on finite candidate sets $\mathcal{S}_{u}^{i}$ but depends directly on $\lambda_{a}$ rather than $\tilde{\lambda}_{a}$. However, as argued in Wang et al. [2023b], bounded-variance assumptions may be unrealistically strong in practice and sometimes conflict with realistic data-generation processes.
Why We Adopt ˜3.1 in the Main Body.
To remain consistent with state-of-the-art clustering of bandits work [Wang et al., 2023a, Dai et al., 2024a, Wang et al., 2025] and to better model realistic offline data (e.g., those generated by logging policies that induce dependencies or violate bounded-variance conditions), we adopt ˜3.1 in the main body. This assumption provides an interpretable and practical balance: it requires finite candidate sets (a natural fit for recommendation systems and ranking tasks) and makes use of the smoothed parameter $\tilde{\lambda}_{a}$, which better captures the effective coverage of the action space under dependent samples. Importantly, our theoretical guarantees remain robust: they also hold under ˜A.1, with only minor substitutions in the analysis. This illustrates the generality of our framework.
Remark A.2 (Beyond Regularity Assumptions).
Finally, we consider scenarios where neither ˜3.1 nor ˜A.1 holds. In our proofs (see Appendix˜C), the purpose of the regularity assumption is to ensure a lower bound on the minimum eigenvalue of the information matrix $M_{u}$ defined in Equation˜1 (i.e. $\lambda_{\min}(M_{u})$), so that the confidence interval $\text{CI}_{u}$ can be controlled in terms of $N_{u}$, the number of samples, as represented in Lemma˜D.2. Without regularity, the denominator of $\text{CI}_{u}$ must instead depend directly on $\sqrt{\lambda_{\min}(M_{u})}$, as shown in Equation˜11. In this case, all performance bounds would explicitly depend on $\lambda_{\min}(M_{u})$ rather than sample size. While this generalization is conceptually straightforward, in order to keep consistent with the standard assumptions used throughout the clustering of bandits literature [Gentile et al., 2014, Li and Zhang, 2018, Li et al., 2019, Liu et al., 2022, Wang et al., 2023a, Dai et al., 2024a, Li et al., 2025, Wang et al., 2025], we still adopt ˜3.1 in the main body and leave detailed analysis of non-regular settings as a promising direction for future work.
Appendix B Additional Theoretical Results
B.1 Fixed-Parameter Form for Theorem˜4.4
By refining the role of $\alpha_{r}$ and $\alpha_{w}$, we can derive an upper bound for Theorem˜4.4 with fixed parameters. Specifically, we replace $N_{\mathcal{V}_{\hat{\gamma}}(u_{\text{test}})}$ with $N_{\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})}$ in the denominator of the first term, select $\alpha_{r}$ at the lower end of its range (to conservatively estimate a subset of $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$), and $\alpha_{w}$ at the upper end (to conservatively estimate a superset of $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$). This yields the concrete form presented in Corollary˜B.1.
Corollary B.1 (Concrete Form of Theorem˜5.2).
Define $\tilde{\mathcal{R}}_{\hat{\gamma}}(u)$ and $\tilde{\mathcal{W}}_{\hat{\gamma}}(u)$ as the concrete instances of $\mathcal{R}_{\hat{\gamma}}(u)$ and $\mathcal{W}_{\hat{\gamma}}(u)$ in Lemma˜4.2, obtained by setting $\alpha_{r}=\frac{\sqrt{\tilde{\lambda}_{a}}}{4(\alpha+1)\sqrt{\log(2U/\delta)\max\{2,d\}}}$ and $\alpha_{w}=\frac{\sqrt{\tilde{\lambda}_{a}}}{2(\alpha-1)\sqrt{\log(2U/\delta)}}$. Let $\tilde{\mathcal{V}}_{\hat{\gamma}}(u)=\tilde{\mathcal{R}}_{\hat{\gamma}}(u)\cup\tilde{\mathcal{W}}_{\hat{\gamma}}(u)$ and define $\eta_{\tilde{\mathcal{W}}_{\hat{\gamma}}(u)}=\frac{|\tilde{\mathcal{W}}_{\hat{\gamma}}(u)|\lambda+N_{\tilde{\mathcal{W}}_{\hat{\gamma}}(u)}}{|\tilde{\mathcal{V}}_{\hat{\gamma}}(u)|\lambda+N_{\tilde{\mathcal{V}}_{\hat{\gamma}}(u)}}.$ Then, with probability at least $1-2\delta$, the suboptimality of Off-C2LUB satisfies
@@ Appendix B Additional Theoretical Results
$\textnormal{SubOpt}(u_{\textnormal{test}},\textnormal{Off-C}^{2}\textnormal{LUB},\mathcal{A}_{\textnormal{test}})\leq\tilde{O}\left(\sqrt{\frac{d}{\tilde{\lambda}_{a}N_{\tilde{\mathcal{R}}_{\hat{\gamma}}(u_{\text{test}})}}}+\frac{\eta_{\tilde{\mathcal{W}}_{\hat{\gamma}}(u_{\text{test}})}\hat{\gamma}}{\tilde{\lambda}_{a}}\right).$ (7)
B.2 Additional Discussions on $\hat{\gamma}$ Selection Policies
Theorem B.2 (Effect of Underestimation Policy).
With $\hat{\gamma}$ chosen according to the underestimation policy in Equation˜5 and $\alpha_{w}^{\prime}=\frac{\sqrt{\tilde{\lambda}_{a}}}{4(\alpha+1)\sqrt{\log(2U/\delta)\max\{2,d\}}}$, any user $v$ in the heterogeneous neighbor set $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ of Lemma˜4.2 also satisfies $\frac{1}{\sqrt{N_{u_{\text{test}}}}}+\frac{1}{\sqrt{N_{v}}}\geq\alpha_{w}^{\prime}\left\|\bm{\theta}_{u_{\text{test}}}-\bm{\theta}_{v}\right\|_{2}.$
Remark B.3 (Interpretation of Underestimation Policy).
Since $\Gamma(u,v)\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}$, all users in $M(u_{\text{test}})$ must belong to clusters different from $u_{\text{test}}$. By taking the minimum over these values, the underestimation policy enforces a small $\hat{\gamma}$, which substantially limits the size of $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$. This prevents the inclusion of highly biased heterogeneous samples, thereby reducing systematic bias in estimation. The additional sample-complexity condition in Theorem˜B.2 further guarantees that only heterogeneous users with insufficient data can enter $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, minimizing their overall impact. However, this conservative choice of $\hat{\gamma}$ also restricts the homogeneous set $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$, leaving fewer samples available for aggregation and thus amplifying noise. The underestimation policy is therefore most suitable for safety-critical applications where bias is more damaging than noise, such as personalized medicine or risk-sensitive decision making.
Theorem B.4 (Effect of Overestimation Policy).
With $\hat{\gamma}$ chosen according to the overestimation policy in Equation˜5, if $M(u_{\text{test}})\neq\emptyset$, then $\hat{\gamma}\geq\gamma$.
Remark B.5 (Interpretation of Overestimation Policy).
By construction, the overestimation policy ensures $\hat{\gamma}\geq\gamma$. This guarantees that more homogeneous users of $u_{\text{test}}$ with sufficient data will be included in $\mathcal{V}_{\hat{\gamma}}(u_{\text{test}})$ under the looser requirement of sample size as represented in Lemma˜4.2. As a result, the aggregated dataset becomes significantly larger, which decreases variance and reduces noise in preference estimation. The drawback, however, is that this looser threshold also admits some heterogeneous users into $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, thereby introducing additional bias. The overestimation policy is thus well-suited for data-rich environments where variance dominates bias—for example, recommender systems with high-dimensional preference vectors, where aggregating large amounts of diverse data is essential for stable learning.
Appendix C Detailed Proofs
C.1 Proof of Lemma˜4.2
@@ Appendix C Detailed Proofs
We first define several events to bound the estimates. Let $\beta(\mathcal{S},\delta)=\sqrt{d\log\left(1+\frac{\sum_{u\in\mathcal{S}}N_{u}}{|\mathcal{S}|\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{|\mathcal{S}|\lambda}$ for any set $\mathcal{S}$. We begin by defining the following confidence event for each user $u$:
@@
By Lemma D.1, we have $\mathbb{P}[\mathcal{F}_{v}^{\prime}(\delta)]\geq 1-\frac{\delta}{2U}$ for each $v\in\mathcal{U}$. Therefore, applying the union bound over all users yields:
@@
Furthermore, for both $\mathcal{R}_{\hat{\gamma}}(u)$ and $\mathcal{W}_{\hat{\gamma}}(u)$, the clustering condition in Equation˜2 requires that a user $v$ can be included in either set only if both $N_{u}$ and $N_{v}$ are at least $\frac{16}{\tilde{\lambda}_{a}^{2}}\log\left(\frac{8dU}{\tilde{\lambda}_{a}^{2}\delta}\right)$. Hence, under the case where $N_{u}<\frac{16}{\tilde{\lambda}_{a}^{2}}\log\left(\frac{8dU}{\tilde{\lambda}_{a}^{2}\delta}\right)$, the sets $\mathcal{R}_{\hat{\gamma}}(u)=\{u\}$ and $\mathcal{W}_{\hat{\gamma}}(u)=\emptyset$ since the cluster condition in Equation˜2 does not hold for any user $v$.
For the non-trivial case where $N_{u}\geq\frac{16}{\tilde{\lambda}_{a}^{2}}\log\left(\frac{8dU}{\tilde{\lambda}_{a}^{2}\delta}\right)$, we apply ˜3.1, Lemma J.1 in Wang et al. [2023a], Lemma 7 in Li and Zhang [2018] and the union bound to obtain that for any $v\in\mathcal{R}_{\hat{\gamma}}(u)\cup\mathcal{W}_{\hat{\gamma}}(u)$, with probability at least $1-\frac{3}{2}\delta$, the following inequality holds:
$\left\|\bm{\theta}_{v}-\hat{\bm{\theta}}_{v}\right\|_{2}\leq\frac{\left\|\bm{\theta}_{v}-\hat{\bm{\theta}}_{v}\right\|_{M_{v}}}{\sqrt{\lambda_{\min}(M_{v})}}\leq\frac{\beta(\{v\},\delta)}{\sqrt{\lambda+\tilde{\lambda}_{a}N_{v}/2}}\leq\text{CI}_{v}.$